OpenStudy (sphott51):

(binominal experiments and simulations) six coins are tossed. what is the probability that the following occurs? 1. 3 heads and 3 tails

satellite73 (satellite73):

3 heads and three tails? formula would be \[\binom{6}{3}\left(\frac{1}{2}\right)^6\]

satellite73 (satellite73):

where \[\binom{6}{3}=\frac{6\times 5\times 4}{3\times 2}\]

OpenStudy (sphott51):

Isn't formula P(r) = C(n,r) P^r q^n-r

OpenStudy (zarkon):

\[C(n,r)= \binom{n}{r}\] \[C(n,r)p^rq^{n-r}\] \[C(6,3)\left(\frac{1}{2}\right)^3\left(\frac{1}{2}\right)^{6-3}\] \[=C(6,3)\left(\frac{1}{2}\right)^3\left(\frac{1}{2}\right)^{3}\] \[=C(6,3)\left(\frac{1}{2}\right)^{3+3}=C(6,3)\left(\frac{1}{2}\right)^{6}\] \[=\binom{6}{3}\left(\frac{1}{2}\right)^{6}\] which is what satellite73 said

OpenStudy (nuke_ur_friend):

@i_am_a_penguin

OpenStudy (nuke_ur_friend):

@timeflybyegirl